Why does this problem end in only two sig figs? 4.18 - 58.16 x (3.38-3.01) Please explain
Answer #1

According to rule of rounding/significant figures, The final answer should be rounded to the least significant figures.

Let's do the calculation here:

3.38-3.01 = 0.37 (2 sig figs.)

Now when we do next step, (58.16 x 0.37 = 21.5192 ), according to significant figure rules, the answer should be rounded to the least significant figures (which here is 2).

Hence, 58.16 x 0.37 = 22

Next step also, (4.18 - 22 = -17.82), according to significant figure​​​​​​​ rules, the answer should be rounded to least significant figures (which here is 2).

So the final answer should be rounded to 2 significant figures.

Hence, 4.18 - 22 = -18